OpenStudy (anonymous):

when dividing 6^-9/6^10 do you just divide the square root of both these?

OpenStudy (anonymous):

6^ -9 / 6^10 = 6^(-9-10) = 6^-19

OpenStudy (anonymous):

you subtract the exponents

OpenStudy (anonymous):

how come you dont do anything with the integers?

OpenStudy (anonymous):

it is 6^-9/6^10 = 6^(-9-10) = 6^(-19)

OpenStudy (anonymous):

you dont because its a the same integer (6) - the law of exponents take a simple example x^3 / x^2 = x^1 or x

OpenStudy (anonymous):

ok just making sure and thanks

OpenStudy (anonymous):

yw